3(x, y) x=2v - Bu, y = 100 + 2v implies a(u, v)
Answer #1

& Solution :- Sunce x=20-84 y=100+20 miplies acxig) auro) on you au olou an too o Suice x=20-84 one 04 -8 a - and ya Lou +20IF YOU FOUND THIS ANSWER USEFUL THEN PLEASE LIKE PLEASE